The load connected between each line and the neutral of a 415 volts, 50hertz 3 phase Circuit consist of a capacitor of 31.8 micro ferad in series with a resistor of 100 ohms. CALCULATE the line current and the Total Power.


1
Expert's answer
2021-03-08T03:08:42-0500

Resistance, "R=100\Omega"

Capacitance, "C=31.8\mu F"

Line Voltage, "V_L=415V"

Frequency of AC supply, "f=50Hz"

Impedance, "Z=\sqrt{R^2+(\dfrac{1}{2\pi fC})^2}"

"\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space=\sqrt{(100)^2+(100.10)^2}"

"\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space=141.50\space\Omega"

"V_{Ph}=\dfrac{V_L}{\sqrt{3}}=\dfrac{415}{\sqrt{3}}=239.60\space V"

Power factor, "cos\phi=\dfrac{R}{Z}=0.706"

"I_{Ph}=\dfrac{V_{Ph}}{Z_{Ph}}=1.69\space A"

Line Current, "I_{L}=I_{Ph}=1.69\space A"

Total Power absorbed , "P_{total}=3\times I_{Ph}^2\times R_{Ph}=3\times 286.72=860.16\space W"